OpenStudy (anonymous):

Find the product of 21 and (6.3 • 10^-4). 0.1323 • 10^-1 1.323 • 10^-3 1.323 • 10^-2 1.323 • 10^2

OpenStudy (mrnood):

first multiply 21 * 6.3 and write it in the exponent form post you answer here

OpenStudy (anonymous):

132.3

OpenStudy (mrnood):

can you write that in 'exponent form ? like 1.56 * 10^2

OpenStudy (anonymous):

Is it 132.3 * 10^0?

OpenStudy (mrnood):

that is one way to write it can you write it as 1.323 * 10^?

OpenStudy (anonymous):

is it possible, you're asking?

OpenStudy (mrnood):

no - it is possible can you do it - what is the value of ? so that 132.3 is the same as 1.323 *10^?

OpenStudy (anonymous):

I know that you have to move the decimal over two spaces to the left by dividing it by 100. That IS correct, right? XD

OpenStudy (mrnood):

yes 132.3= 1.323*100 = 1.323 *10^? what is the value of ? in the above?

OpenStudy (anonymous):

I have no clue. 1.323*100 is 132.3 but that's what we started with...

OpenStudy (mrnood):

the point of the exercise is to be able to express numbers in exponent form (or sometimes called scientific notation) you can see maybe that (for example) 1101 = 110.1 *10^1 1101 = 11.01 * 10^2 1101 = 1.101 * 10^3 1101 = .1101 * 10 ^4 SO using the same idea can you look at 132.3= 1.323*100 = 1.323 *10^?

OpenStudy (anonymous):

is it... 1.323 • 10^-3

OpenStudy (anonymous):

is it... 1.323 • 10^-3

OpenStudy (mrnood):

132.3= 1.323*100 we are multiplying by 100 how would you write 100 = 10^?

OpenStudy (anonymous):

eh... I dont know

OpenStudy (mrnood):

Well I'm afaid you are probably not ready to tackle questions like this then. The key point is that we can move the decimal point where we want - as LONG AS we multiply by 10^? to compensate. BUT you must know that 10^0 =1 10^1 = 10 10^2 = 100 10^3 = 1000 etc.

OpenStudy (anonymous):

Oh! Oh that makes sense

OpenStudy (mrnood):

how would you write 100 = 10^?

OpenStudy (anonymous):

100 would be 10^2

OpenStudy (mrnood):

OK so lets go back toi the original question 21 *6.3 • 10^-4 you worked out was the same as 132.3 *10^4 which is the same as 1.323 *10^2 * 10^-4 so what happens when you multiply 2 numbers with exponents? 10^2 * 10^-4 = 10 ^?

OpenStudy (anonymous):

When you multiply 2 numbers with exponents dont you add the exponents and if the bases are the same then you keep them as that number? So this would be 10^-2?

OpenStudy (mrnood):

excellent - so you have your answer now

OpenStudy (anonymous):

Ah! Thank you so much